OpenStudy (anonymous):

I will fan and give a medal to someone who helps me with 5 math questions :)

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

rewrite the radical as a rational exponent \[\sqrt[4]{7^{5}}\] A. \[7\frac{ 5 }{ 4 }\] B. \[7^{20}\] C. 7 D. \[7\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

2.) Explain how the Quotient Of Powers was used to simplify this expression \[\frac{ 2^{5} }{ 8 } = 2^{2}\] A. By finding the quotient of the bases to be, 1/4 and cancelling common factors B. By finding the quotient of the bases to be, 1/4 and simplifying the expression C. By simplifying 8 to \[2^{3}\] to make both power bases two, and subtracting the exponents D. By simplifying 8 to \[2^{3}\] to make both power bases two, and adding the exponents

OpenStudy (anonymous):

@gerryliyana ^

OpenStudy (anonymous):

C. By simplifying 8 to 2^3 to make both power bases two, and subtracting the exponents

OpenStudy (anonymous):

\[\frac{ 2^5 }{ 8 }=\frac{ 2^5 }{ 2^3 }=2^{5-3} = 2^2\]

OpenStudy (anonymous):

3) Rewrite the rational exponent as a radical exponent \[\left(\begin{matrix}3\frac{ 2 }{ 3 } \\ \end{matrix}\right) \frac{ 1 }{ 6 }\] A. \[\sqrt[6]{3}\] B. \[\sqrt[9]{3}\] C. \[\sqrt[18]{3}\] D. \[\sqrt[6]{3^{3}}\]

OpenStudy (anonymous):

@gerryliyana ^

OpenStudy (anonymous):

you mean \(3^{\frac{ 2 }{ 3 }}\) or \(3\frac{ 2 }{ 3 }\) ?

OpenStudy (anonymous):

\[3\left(\begin{matrix}2 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

the first or the second ??

OpenStudy (anonymous):

second

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

@theopenstudyowl help ^^^^^^

OpenStudy (anonymous):

i am back

OpenStudy (anonymous):

sorry, i was busy eating :)

OpenStudy (anonymous):

that's fine :)

OpenStudy (anonymous):

i just want to get these questions answered so i can go take my shower and get ready to go with my sis and my mom

OpenStudy (anonymous):

\[\left( 3^{\frac{ 2 }{ 3 }} \right)\frac{ 1 }{ 6 } = 3^{\frac{ 2 }{ 3 }}3^{-\frac{ 3 }{ 2 }}=3^{\frac{ 2 }{ 3 }-\frac{ 3 }{ 2 }}=3^{\frac{ 2 \times 2 -3 \times 3}{ 6 }}=3^{\frac{ 4 -9}{ 6 }}=3^{\frac{ -5 }{ 6 }}\]

OpenStudy (anonymous):

sorry i was wrong

OpenStudy (anonymous):

i am gonna go with b?

OpenStudy (anonymous):

i don't know if it's a, b, c, or d

OpenStudy (anonymous):

what do you mean you wrong?

OpenStudy (anonymous):

my last work was wrong

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

it is \((3^{\frac{ 2 }{ 3 }})\), not \(3\left( \frac{ 2 }{ 3 } \right) \)

OpenStudy (anonymous):

oh. ok

OpenStudy (anonymous):

you should transform \(6^{-1}\ to\ 3^{-x}\). You should find x

OpenStudy (anonymous):

and that would give me b?

OpenStudy (anonymous):

where did you get b?

OpenStudy (anonymous):

you said to transform \[\frac{ 1 }{ 6}\] to \[\frac{ x }{ 3 }\] and i multiplied 3 x 1 and 3 x 6 which would give me.............18

OpenStudy (anonymous):

i'm wrong aren't I?

OpenStudy (anonymous):

Nooo.,

OpenStudy (anonymous):

It is not what I mean

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 } = 3^x\] what is "x" ?

OpenStudy (anonymous):

6 or 9? i am really bad at this X(

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

is it wrong?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aw man :(

OpenStudy (anonymous):

i thought i got it right

OpenStudy (anonymous):

hmm i think you wrong

OpenStudy (anonymous):

its should be \[3^{\left( {\frac{ 2 }{ 3 }} \right)\frac{ 1 }{ 6 }}\]

OpenStudy (anonymous):

So

OpenStudy (anonymous):

\[3^{\left( \frac{ 2 }{ 3 } \right)\frac{ 1 }{ 6 }}=3^{\frac{ 2 }{ 18 }}=3^{\frac{ 1 }{ 9 }}\]

OpenStudy (anonymous):

the answer is b

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

you made a typo man

OpenStudy (anonymous):

you know what i mean?

OpenStudy (anonymous):

the answer is B

OpenStudy (anonymous):

next question, please

OpenStudy (anonymous):

Rewrite the rational exponent as a radical by extending the properties of integer exponents \[2\frac{ 3 }{ 4 } \over 2\frac{ 1 }{ 2 }\] A. \[\sqrt[8]{2^{3}}\] B. \[\sqrt{2\frac{ 3 }{ 4 }}\] C. \[\sqrt[4]{2}\] D. \[\sqrt{2}\]

OpenStudy (anonymous):

\[\frac{ 2^{\frac{ 3 }{ 4 }} }{ 2^{\frac{ 1 }{ 2 }} }=2^{\frac{ 3 }{ 4 }-\frac{ 1 }{ 2 }}= 2 ^{\frac{ 1 \times 3 - 2 \times 1}{ 4 }} = 2^{\frac{ 3-2 }{ 4 }}=2^{\frac{ 1 }{ 4 }} =\sqrt[4]{2}\]

OpenStudy (anonymous):

that's all

OpenStudy (anonymous):

Now, you know what should you choose

OpenStudy (anonymous):

C

OpenStudy (anonymous):

you got the point man!

OpenStudy (anonymous):

can you help me with one more question?

OpenStudy (anonymous):

ok, but don't forget to fan me and take the time to write a testimonial about me, i will appreciate it!

OpenStudy (anonymous):

A rectangle has a length of \[\sqrt[3]{81}\] and a width of \[3\frac{ 2 }{ 3 }\]. Find the area of the rectangle. A. \[3\frac{ 2 }{ 3 }\] inches squared B. \[3\frac{ 8 }{ 3 }\] inches squared C. 9 inches squared D. \[9\frac{ 2 }{ 3 }\] inches squared

OpenStudy (anonymous):

Help? @gerryliyana

OpenStudy (anonymous):

@TheSmartOne quick help? ^

OpenStudy (anonymous):

\[\sqrt[3]{81}= 81^{\frac{ 1 }{ 3}}=\left( 3^4 \right)^{\frac{ 1 }{ 3 }}=3^{\frac{ 4 }{ 3 }}= (3^2)^{\frac{ 2 }{ 3 }}=9^{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

Extra quick help from me., are you satisfied?

OpenStudy (anonymous):

@rikkibracamonte1999 are you still here?